OpenStudy (anonymous):

A rectangular prism is truncated removing 20% of its height. The original height of the prism was 100 inches and it has a length of 5 inches and a width of 10 inches. What is the volume of the truncated prism? 4,000 in3 2,000 in3 3,000 in3 5,000 in3

OpenStudy (dumbcow):

20% of 100 = 20 100-20 = 80 (new height) volume= 80*5*10

OpenStudy (anonymous):

thanks